# Increasing Efficiency and Result Reliability of Continuous Benchmarking for FaaS Applications

### v1

@@ Abstract.
In this paper, we propose DuetFaaS, an approach which adapts duet benchmarking to FaaS applications. With DuetFaaS, we deploy two versions of FaaS function in a single cloud function instance and execute them in parallel to reduce the impact of platform variability. We evaluate our approach against state-of-the-art approaches, running on AWS Lambda. Overall, DuetFaaS requires fewer invocations to accurately detect performance regressions than other state-of-the-art approaches. In 99.65% of evaluated cases, our approach provides smaller confidence interval sizes than the comparing approaches, and can reduce the size by up to 98.23%.
@@ 1. Introduction
The naive approach to continuous FaaS benchmarking, i.e., deploying two versions of a function and iteratively invoking both, requires hundreds to thousands of function invocations in order to yield meaningful results (Bermbach et al., 2017; Leitner and Cito, 2016; Bermbach, 2017a; Schirmer et al., 2023). This is simply a result of the high level of abstraction of FaaS platforms, which build on complex hardware and software stacks, and are thus subject to performance variability (Scheuner et al., 2022; Scheuner, 2022). Previous research has proposed methodologies and strategies to reduce and mitigate the impact of such influences (Grambow et al., 2023a, 2019, 2021). Thus far, all strategies use either parallel, independent function invocations, which can be affected by performance variability between the machines the different function instances are deployed on, or sequential function invocations within the same function instances, which is subject to temporal performance variability. We instead propose adapting the concept of duet benchmarking (Bulej et al., 2020) from microbenchmarks to FaaS applications: both function versions are deployed in a shared cloud function instance and are executed in parallel, each running on a separate CPU core but with shared memory. This minimizes influence of temporal performance variation and heterogeneous underlying hardware. In this regard, we make the following contributions:
@@
Through our experiments, we show that DuetFaaS provides more accurate results in 99.65% of experiments and requires as few as 100 function invocations to receive results with a confidence interval with a width of less than 2 percent of the median. Compared to the state of the art, this means that DuetFaaS provides almost always more stable results with significantly less repetitions, i.e., significantly faster and less expensive.
@@ 2.3. Continuous FaaS Application Benchmarking
Duet benchmarking has however not been implemented for FaaS application releases to the best of our knowledge. Duet benchmarking aims to guarantee resource fairness for both function versions by executing them in parallel on the same instance. The artifacts are usually containerized and provided the same amount of resources (CPU, memory, etc.), which also causes occurring uncontrollable factors to influence both versions equally (Bulej et al., 2020). This process enables a fair benchmarking procedure when wanting to efficiently detect performance variation in software application releases. Duet benchmarking was originally introduced to improve measurement accuracy on cloud instances, but has been extended to other fields (Bulej et al., 2020). More recently it was shown that it is well suited for microbenchmarks, and early findings also showed accuracy improvements with benchmarking suites (Bulej et al., 2020, 2019).
@@ 3.1. Parallel Isolated Execution
The amount of factors that we need to control exceeds that of a native deployment, hence is why we made the choice to deploy a web server of sorts as a FaaS function. This allows us to pass configuration options to the server, which ensure both the parallel execution and the distribution of artifacts across separate CPU cores. We can use the primary process running on our server to coordinate worker processes, which can then be told to run each artifact in parallel. This process is also responsible for isolating workers on separate cores, which provides one of the requirements for duet benchmarking. It also passes one release to each core, effectively isolating them. We can also use the same primary process directly as the scheduler. Once both worker processes are ready, we can coordinate the parallel execution of function artifacts by synchronously starting the artifacts on the workers.
@@ 3.3. Pipeline Integration
As our function is deployable to a public cloud platform, we can also use an action within a deployment pipeline to run fully automated experiments. We can set up the action to monitor the version control system for relevant changes, which then trigger an automatic deployment, benchmark execution, and destruction of duet benchmark experiments. This is also another reason why a local deployment would not suffice, as it would prove difficult to make every possible local environment with compatible with a deployment pipeline.
@@ 4.1. Proof-of-Concept Prototype
We build our prototype in AWS and rely on a repository built specifically for AWS111https://github.com/awslabs/aws-lambda-web-adapter. This allows us to deploy a web server to Lambda, which can host our function artifacts. We can configure this FaaS function to run an Express server in cluster mode. This creates a master process, which creates two worker processes with each of the workers being run on a separate core. Once each worker is ready, we use the master process to synchronously call the function executions and enable duet benchmarking within an actual FaaS environment.
@@ 4.2. Experiment Design
Our experiments with DuetFaaS aim to show to what extent our approach can improve the accuracy of detecting performance regressions in FaaS functions. To this end, we deploy two versions of a FaaS function, execute performance benchmarks against that function using different benchmarking approaches, and then compute the size of the bootstrapped results’ confidence intervals. Additionally, we are interested in how many benchmark executions are necessary to achieve that confidence interval. Ideally, we reduce both the confidence interval size (meaning results are more accurate) while also reducing the number of necessary executions (meaning the performance evaluation is cheaper and faster to execute). Naturally, there is a trade-off between measuring more data for a smaller confidence interval and reducing the number of benchmark execution, which is why we measure confidence interval size with a varying number of benchmark executions included in the result computation, evaluating the compared approaches along this trade-off. We deploy two functions to AWS Lambda, a CPU-intensive neural network mutation (CPU) (Susnic, 2017) and a memory-intensive implementation of the Sieve of Eratosthenes algorithm used to calculate prime numbers (MEM) (Rev. Samuel Horsley, 1772). Both functions are implemented in JavaScript. Each function is allocated 512MB of memory, which yields roughly 0.289 vCPUs (Services, 2021). In our DuetFaaS approach, we allocate 1024MB of memory to the function (0.58 vCPUs) as two instances are invoked in parallel.
@@
We perform A/A and A/B benchmarks of each function. For A/A benchmarks, we ‘compare’ two identical implementations of the function, which ideally yields no detected performance regression. In our A/B benchmarks, we artificially inject a 5% performance regression in each function by modifying the code accordingly, i.e., performing 5% more mutation in the CPU function and calculating 5% more prime numbers in the MEM function.
@@
We compare the DuetFaaS approach against an improved traditional independently called benchmark and sequential randomized (RMIT / faasterBench) invocation. In the improved independent approach, we invoke each function version individually. However, the versions are not deployed individually, and instead are put in a wrapper artifact. This leads to the executing instance being the same, which would very likely not occur if we deployed both versions independently. Sequential randomized invocation is adapted from Grambow et al. (Grambow et al., 2023a). Here, both function versions are deployed together and invoked sequentially within the same function executor. A wrapper decides the execution order for each invocation, implementing an RMIT strategy. We run 1,500 benchmark repetitions for each approach, calculate the median performance change for each benchmark, and calculate a confidence interval with a confidence level of 99% using bootstrapping for the performance changes. To assert reproducibility, we repeated all experiments three times across multiple times-of-day (Schirmer et al., 2023) and observed very similar results. Further, we note that we remove cold start executions from our results.
@@ 4.3. Performance Regression Detection
We will focus on one of our experiment runs in this analysis, however the other runs produced similar results.
In Figure 4, we show the performance regression that each approach detects with 1,500 made calls, i.e, from one of our experiment runs. The other runs produced similar results.
@@
Across all four experiments, duet benchmarking’s average CI size is only 0.28 percentage points when including 1,500 benchmark experiments, which is 90% / 87.94% smaller than both the traditional independent invocation (2.80) and the randomized sequential one (2.32). Therefore, we recommend the duet benchmarking methodology for benchmarking FaaS applications when running longer experiments.
We find the CPU function to provide significant improvements over currently used approaches. Both configurations provided an interval size of 0.0 percentage points when using duet benchmarking, compared to the 1.96 of randomized sequential and 2.47 of traditional independent. This highlights the staggeringly high level of accuracy that duet benchmarking is able to provide, even in FaaS environments.
@@ 4.4. Impact of Sample Size
Our experiments show that a similar level of interval size is already reached with as few as 100 results (depending on the configuration), with all scenarios requiring less than 200 calls made to achieve such a value. This is noticeably different to the other approaches, which require more calls before a stable accuracy level is achieved. When only including 100 results, duet benchmarking already provides an average interval size of only 0.36 percentage points, while the traditional independent provides a size of 6.31. The randomized sequential methodology is even worse at this point, with an interval of over 7.40 percentage points. We notice large improvements from both comparing approaches however when including more results, which are still far from the level of accuracy that duet benchmarking is able to provide with fewer results. Worthy of note is that duet benchmarking is commonly able to provide more accurate results with fewer repetitions than comparing results with more. Considering the noticeable accuracy stagnation in the traditional and randomized sequential methodology leads us to conclude that these approaches would never reach the level of accuracy that duet benchmarking is able to achieve with only 100 results.
@@ 4.5. Key Findings
Duet benchmarking provides more accurate intervals than both the traditional independent and randomized sequential approach by a significant margin. We find that applying duet benchmarking can reduce interval size by up to 98.23% depending on the experiment configuration. We also find that duet benchmarking works well for varying function types and detects the injected performance regression accurately.
Additionally, duet benchmarking requires fewer results to provide an accurate and stable confidence interval. While other approaches require up to 700 results to provide a stable confidence interval size, duet benchmarking needs only slightly more than 100, again depending on the experiment.
Overall, duet benchmarking provides better interval sizes than the other approaches in 99.65% of evaluated cases, and provides smaller sizes with fewer results than other approaches with more repetitions.
@@ 5. Discussion
Our experiments show that duet benchmarking provides accurate results with as few as 100 results and does not necessarily require more invocations, as a high degree of result accuracy is already reached with that many repetitions. This indicates that a short benchmarking step is already enough to provide a usable result for the FaaS application, which can help speed up the deployment process.
While we consider duet benchmarking to already be usable for specific FaaS applications, further improvements and implementations must be made to extend duet benchmarking to other FaaS environments.
@@
Our results however showed that our approach works with both types, so we do not suspect that the results would differ heavily.
We plan to include more types of functions to provide a broader comparison ground and further validate our approach.
@@ 6. Conclusion
Our experiments with a proof-of-concept prototype on AWS Lambda showed that duet benchmarking improves benchmark accuracy when compared to the traditional and randomized sequential approach, and provided more narrow confidence intervals with fewer results than other approaches in 99.65% of evaluated cases. When including only 100 results, duet benchmarking provided intervals only 5.7% as large as the other approaches, which helps developers better evaluate their function artifacts quickly. We propose to include duet benchmarking for FaaS application releases, and plan to continue our work to improve the approach.

### v2

@@ Abstract.
In this paper, we propose DuetFaaS, an approach which adapts duet benchmarking to FaaS applications. With DuetFaaS, we deploy two versions of FaaS function in a single cloud function instance and execute them in parallel to reduce the impact of platform variability. We evaluate our approach against state-of-the-art approaches, running on AWS Lambda. Overall, DuetFaaS requires fewer invocations to accurately detect performance regressions than other state-of-the-art approaches. In of evaluated cases, our approach provides equal or smaller confidence interval size. DuetFaaS achieves an interval size reduction in of all evaluated sample sizes when compared to the competitive approaches.
@@ 1. Introduction
The naive approach to continuous FaaS benchmarking, i.e., deploying two versions of a function and iteratively invoking both, requires hundreds to thousands of function invocations in order to yield meaningful results (Bermbach et al., 2017; Leitner and Cito, 2016; Bermbach, 2017a; Schirmer et al., 2023). This is simply a result of the high level of abstraction of FaaS platforms, which build on complex hardware and software stacks, and are thus subject to performance variability (Scheuner et al., 2022; Scheuner, 2022).
Previous research has found that variability differs across FaaS platforms, with performance differences up to being detected (Schirmer et al., 2023). It also has proposed methodologies and strategies to reduce and mitigate the impact of such influences (Grambow et al., 2023a, 2019, 2021). Thus far, all strategies use either parallel, independent function invocations, which can be affected by performance variability between the machines the different function instances are deployed on, or sequential function invocations within the same function instances, which is subject to temporal performance variability. We instead propose adapting the concept of duet benchmarking (Bulej et al., 2020) from microbenchmarks to FaaS applications: both function versions are deployed in a shared cloud function instance and are executed in parallel, with each function being provided the same amount of resources. This minimizes influence of temporal performance variation and heterogeneous underlying hardware. In this regard, we make the following contributions:
@@
Through our experiments, we show that DuetFaaS provides equal or smaller confidence interval sizes in of experiments and requires as few as 100 function invocations to receive results with a confidence interval with a width of less than 2 percent of the median. Compared to the state of the art, this means that DuetFaaS provides almost always more stable results with less repetitions, i.e., is faster and less expensive.
@@ 2.3. Continuous FaaS Application Benchmarking
Duet benchmarking has however not been implemented for FaaS application releases to the best of our knowledge. Duet benchmarking aims to guarantee resource fairness for both function versions by executing them in parallel on the same instance. The artifacts are usually containerized, provided the same amount of resources (CPU, memory, etc.), and then executed in parallel, which also causes occurring uncontrollable factors to influence both versions equally (Bulej et al., 2020). This process enables a fair benchmarking procedure when wanting to efficiently detect performance variation in software application releases. Duet benchmarking was originally introduced to improve measurement accuracy on cloud instances, but has been extended to other fields (Bulej et al., 2020). More recently it was shown that it is well suited for microbenchmarks, and early findings also showed accuracy improvements with benchmarking suites (Bulej et al., 2020, 2019).
@@ 3.1. Parallel Isolated Execution
The amount of factors that we need to control exceeds that of a native deployment, hence is why we made the choice to deploy a web server of sorts as a FaaS function. This allows us to pass configuration options to the server, which ensure both the parallel execution of both artifacts, and allowing us to isolate each artifact to a separate core. We can use the primary process running on our server to coordinate worker processes, which can then be told to run each artifact in parallel. We can also use the same primary process directly as the scheduler. Once both worker processes are ready, we can coordinate the parallel execution of function artifacts by synchronously starting the artifacts on the workers.
@@ 3.3. Pipeline Integration
As our function is deployable to a public cloud platform, we can also use an action within a deployment pipeline to run fully automated experiments. We can set up the action to monitor the version control system for relevant changes, which then trigger an automatic deployment, benchmark execution, and destruction of duet benchmark experiments. This is also another reason why a local deployment would not suffice, as it would prove difficult to make every possible local environment compatible with a deployment pipeline.
@@ 4.1. Proof-of-Concept Prototype
We build our prototype in AWS and rely on a repository built specifically for AWS111https://github.com/awslabs/aws-lambda-web-adapter. This allows us to deploy a web server to Lambda, which can host our function artifacts. We can configure this FaaS function to run an Express server in cluster mode. This creates a master process, which creates two worker processes with each of the workers being run as a separate execution thread. Once each worker is ready, we use the master process to synchronously call the function executions and enable duet benchmarking within an actual FaaS environment.
@@ 4.2. Experiment Design
Our experiments with DuetFaaS aim to show to what extent our approach can improve the accuracy of detecting performance regressions in FaaS functions. To this end, we deploy two versions of a FaaS function, execute performance benchmarks against that function using different benchmarking approaches, and then compute the size of the bootstrapped results’ confidence intervals. Additionally, we are interested in how many benchmark executions are necessary to achieve that confidence interval. Ideally, we reduce both the confidence interval size (meaning results are more accurate) while also reducing the number of necessary executions (meaning the performance evaluation is cheaper and faster to execute). Naturally, there is a trade-off between measuring more data for a smaller confidence interval and reducing the number of benchmark execution, which is why we measure confidence interval size with a varying number of benchmark executions included in the result computation, evaluating the compared approaches along this trade-off. We deploy two functions to AWS Lambda, a CPU-intensive neural network mutation (CPU) (Susnic, 2017) and a memory-intensive implementation of the Sieve of Eratosthenes algorithm used to calculate prime numbers (MEM) (Rev. Samuel Horsley, 1772). Both functions are implemented in JavaScript. Each function is allocated 2048 MB of memory for the traditional and randomized sequential approach. In our DuetFaaS approach, we allocate 3008 MB of memory to the function, as we need to run two artifacts in parallel. This amount of memory allocates at least two cores to our Lambda function, although less would have sufficed (Services, 2021).
@@
We perform A/A and A/B benchmarks of each function. For A/A benchmarks, we ‘compare’ two identical implementations of the function, which ideally yields no detected performance regression. In our A/B benchmarks, we artificially inject an artificial performance regression in each function by modifying the code accordingly, i.e., performing more mutations in the CPU function and increasing the prime number upper threshold by in the MEM function.
@@
We compare the DuetFaaS approach against an improved traditional independently called benchmark and sequential randomized (RMIT / faasterBench) invocation. In the improved independent approach, we invoke each function version individually without the other version. However, the versions are not deployed individually, and instead are put in a wrapper artifact. This leads to the executing instance being the same, which would very likely not occur if we deployed both versions completely independent of one another. Sequential randomized invocation is adapted from Grambow et al. (Grambow et al., 2023a). Here, both function versions are deployed together and invoked sequentially within the same function executor. A wrapper decides the execution order for each invocation, implementing an RMIT strategy. This is then repeated for multiple iterations, with the order being randomly determined each iteration. We run 1,500 benchmark repetitions for each approach, calculate the median performance change for each benchmark, and calculate a confidence interval with a confidence level of using bootstrapping for the performance changes. To assert reproducibility, we repeated all experiments three times across multiple times-of-day (Schirmer et al., 2023) and observed very similar results. Further, we note that we remove cold start executions from our results.
@@ 4.3. Performance Regression Detection
We will focus on one of our experiment runs in this analysis, however the other runs produced similar results. In Figure 4, we show the performance regression that each approach detects with 1,500 made calls, i.e, from one of our experiment runs.
@@
Across all four experiments, duet benchmarking’s average CI size is 0.01 percentage points when including 1,500 benchmark experiments, which is / smaller than both the traditional independent invocation (1.24) and the randomized sequential one (0.19). Therefore, we recommend the duet benchmarking methodology for benchmarking FaaS applications when running longer experiments.
We find the CPU function to provide significant improvements over currently used approaches. Both configurations provided an interval size of 0.0 percentage points when using duet benchmarking, compared to the 0.15 of randomized sequential and 1.40 of traditional independent. This highlights the staggeringly high level of accuracy that duet benchmarking is able to provide, even in FaaS environments.
@@ 4.4. Impact of Sample Size
Our experiments show that a similar level of interval size is already reached with as few as 100 results (depending on the configuration), with all scenarios requiring less than 200 calls made to achieve such a value. This is noticeably different to the other approaches, which require more calls before a stable accuracy level is achieved. When only including 100 results, duet benchmarking already provides an average interval size of only 0.26 percentage points, while the traditional independent provides a size of 1.29. The randomized sequential methodology is better than the traditional at this point, with an interval of over 0.28 percentage points. We notice large improvements by both the randomized sequential and duet approaches compared to the traditional approach. Of note is duet benchmarking’s reduced interval size over both approaches in A/B configurations. Worthy of note is that duet benchmarking is commonly able to provide narrower CIs with fewer repetitions than comparing results with more. Considering the noticeable accuracy stagnation in the traditional and randomized sequential methodology leads us to conclude that these approaches would likely never reach the level of accuracy that duet benchmarking is able to achieve with only 100 results, depending on the configuration.
@@ 4.5. Key Findings
Duet benchmarking provides narrower confidence intervals than both the traditional independent and randomized sequential approach by a significant margin. We find that applying duet benchmarking can reduce interval size by up to depending on the experiment configuration, along with working well for varying function types and being able to detect the injected performance regression accurately. Additionally, duet benchmarking requires fewer results to provide an accurate and stable confidence interval. While other approaches require up to 500 results to provide a stable confidence interval size, duet benchmarking only needs 100.
Overall, duet benchmarking provides equal or better interval sizes than the other approaches in of evaluated cases, and provides smaller sizes with fewer results than other approaches with more repetitions. We propose to include duet benchmarking as a continuous benchmarking step in deployment pipelines.
@@ 5. Discussion
Our experiments show that duet benchmarking provides accurate results with as few as 100 results and does not necessarily require more invocations, as a high degree of result accuracy is already reached with that many repetitions. This indicates that a short benchmarking step is already enough to provide a usable result for the FaaS application, which can help speed up the deployment process. While we consider duet benchmarking to already be usable for specific FaaS applications, further improvements and implementations must be made to extend duet benchmarking to other FaaS environments.
@@
Our results however showed that our approach works with both types, so we do not suspect that the results would differ heavily. Preliminary experiments have been run with more complex functions, and similar results as presented in this paper were achieved. We aim to publish these in future work, along with further extending the implemented function types to provide a broader comparison ground and further validate our approach.
@@ 6. Conclusion
Our experiments with a proof-of-concept prototype on AWS Lambda showed that duet benchmarking improves benchmark accuracy when compared to the traditional and randomized sequential approach, and provided equal or more narrow confidence intervals with fewer results than other approaches in of evaluated cases. When including only 100 results, duet benchmarking reduces the interval size by up to , which helps developers better evaluate their function artifacts quickly. We propose to include duet benchmarking for FaaS application releases, and plan to continue our work to improve the approach.